# Accelerating Reionization Constraints: An ANN-Emulator Framework for the SCRIPT Semi-numerical Model

### v1

@@ 2.3 Parameter constraints using conventional MCMC
For all observational constraints, except (ii) and (iii), we use a multidimensional Gaussian likelihood of the form
@@
where the index $i$ runs over all the data points used. For data with asymmetric error bars, we use the upper uncertainty when the model estimate lies above the data point, and the lower uncertainty when it lies below. For constraints (ii) and (iii), where we have limits instead of measurements, we use a likelihood given by [gharaConstrainingIntergalacticMedium2020, maityConstrainingReionizationThermal2022]
@@
To sample the parameter space and generate the posterior probability distributions, we use the publicly available package Cobaya [torradoCobayaBayesianAnalysis2019a, torradoCobayaCodeBayesian2021a]. Specifically, we use its implementation of the Metropolis-Hastings sampler [metropolisEquationStateCalculations1953, lewisEfficientSamplingFast2013a]. The samples are drawn using 8 parallel chains [lewisCosmologicalParametersCMB2002a, lewisEfficientSamplingFast2013a]. The chains are assumed to be converged when the Gelman-Rubin R1R-1 statistic [gelmanInferenceIterativeSimulation1992] falls below 0.01 for two consecutive evaluations. Once converged, we discard the first 30% of the samples as burn-in and use the rest for further analysis. The resultant posterior distributions are analyzed using the GetDist [lewisGetDistPythonPackage2019] package. We use simulations with two different resolutions to constrain the free parameters, one with a grid size of $\Delta x=8~h^{-1}$ cMpc ($32^{3}$ grid cells), and another with a grid size of $\Delta x=4~h^{-1}$ cMpc ($64^{3}$ grid cells). The resultant parameter constraints for the two resolutions are shown in red in figures 4 and 5. The significant computational cost of this full MCMC run, which is quantified later (see table 4), provides the primary motivation for developing the efficient ANN-based emulator framework described in the next section.
@@ 3 ANN-Based Emulator
In this section, we present our ANN-based emulator approach, beginning with a brief overview of ANNs, followed by a detailed description of how they are used to construct emulators of $\chi^{2}(\theta)$, including dataset generation, training methodology, and integration within an MCMC framework. The same observational constraints and priors used in the previous section are adopted to enable a direct comparison between the two approaches. All the emulators discussed here are implemented using the TensorFlow Keras API [tensorflow2015-whitepaper, chollet2015keras].
@@ 3.1 A brief on ANNs
by iteratively updating the weights and biases of the network over multiple training cycles, called epochs, with updates occurring after processing mini-batches of data. The training uses an algorithm called backpropagation [rumelhartLearningRepresentationsBackpropagating1986], which applies the chain rule to compute gradients of the loss function with respect to the weights and biases. These gradients are then used by an optimization algorithm to update the network’s weights and biases. The step size of these updates is controlled by the learning rate, which determines how quickly the model adjusts its weights during training.
@@ 3.2 General methodology for emulator training and MCMC integration
Building upon the basic principles outlined above, our primary objective is to train an ANN-based emulator that can accurately and rapidly interpolate the chi-squared value, $\chi^{2}(\theta)$, for any given parameter vector $\theta$ within the prior volume. In particular, the emulator must interpolate accurately in the high-likelihood (low $\chi^{2}$) region, where precise evaluations of $\chi^{2}(\theta)$ are most critical for parameter inference. This trained emulator can then be embedded within a conventional MCMC framework to efficiently explore the parameter space. We choose to emulate $\chi^{2}(\theta)$ directly rather than the individual observables entering the likelihood. This approach significantly simplifies the learning task: $\chi^{2}$ is expected to be a smooth scalar function of the parameters, avoiding the challenges of jointly emulating several observables with different dynamic ranges and noise properties. This choice has also proven robust in earlier emulator frameworks for SCRIPT [maityFastMethodReionization2023, choudhuryGPRbasedEmulatorSeminumerical2024], and it provides a scalable strategy for higher-dimensional extensions of the model. Moreover, building an emulator for a single scalar quantity such as $\chi^{2}$ allows us to use a considerably simpler ANN architecture and, consequently, achieve significantly faster training. This section outlines the general framework we use for training, validating, and deploying these emulators.
@@
The emulator is implemented as a multilayer perceptron (MLP) with Rectified Linear Unit (ReLU) activations in all hidden layers. Optimization is performed using the AdamW optimizer [loshchilovDecoupledWeightDecay2019] with a weight decay factor of 0.004, and all layer weights are initialized using the Glorot scheme [glorotUnderstandingDifficultyTraining2010]. A given training dataset (described in section 3.3) is partitioned into training, model-validation, and test subsets in the ratio 6:2:2, and the batch size is chosen as the power of 2 closest to the square root of the training set’s size.
@@
During each training stage, a grid of candidate networks are trained on the training set, systematically varying the number of layers, the number of neurons per layer, and the learning rate, with the explored ranges listed in table 2. To mitigate the effect of unfavorable weight initialization, which can sometimes yield poor results even for otherwise well-chosen hyperparameters, each candidate network is trained twice using different random seeds. Training continues until the MSE on the model-validation set fails to decrease for 50 consecutive epochs, implementing early stopping to prevent overfitting. Once all candidate networks have been trained, they are evaluated on the test set, and the best-performing emulator is chosen as the trained network with the largest value of
@@ 3.3.1 Results using Latin hypercube sampling
Figure 1: Parameter constraints from the full high-resolution ($N_{\text{grid}}=32$) MCMC run (red) and the ANN-based MCMC runs with the emulator trained on $10^{4}$ parameter vectors sampled using LH sampling from the full prior, and with the priors reduced by factors of 2 and 4. The dashed lines mark the prior boundaries.
@@ 3.3.2 Training set using targeted sampling
In this work, we instead follow the targeted sampling strategy introduced in [maityFastMethodReionization2023], which takes advantage of two features of our setup: (i) the large-scale properties of our model are convergent with respect to simulation resolution, and (ii) full MCMC runs with coarse-resolution simulations are computationally inexpensive. This allows us to sample parameter vectors for the training dataset directly from the MCMC chain of a full MCMC run using inexpensive coarse-resolution simulations, ensuring that samples are concentrated in the high-likelihood regions of parameter space.
@@
In addition, we define a fixed validation set, $\{\theta\}_{\mathrm{val}}$, at the outset. This set is used exclusively to evaluate the convergence of the adaptive sampling procedure and is distinct from the ‘model-validation’ set described in section 3.2, which is used during network training for early stopping. This set is constructed by drawing samples with probability proportional to their weights, ensuring that it provides an unweighted sample representative of the full posterior distribution. Crucially, this set is never used during training, as doing so would cause data leakage, where information from the validation set inadvertently influences training, resulting in overly optimistic performance estimates and poor generalization to unseen parameters. After each training stage, the emulator predicts $\chi^{2}$ for the validation set. The training dataset is then augmented with 300 additional samples, a new emulator is retrained, and the validation predictions are updated. To measure convergence, we compute the KL divergence between successive validation predictions,
@@ 3.4 Parameter constraints using ANN-based MCMC
For our setup, convergence of the adaptive sampling procedure required a training dataset size of 1200 samples for the $N_{\text{grid}}=32$ emulator, and 900 for the $N_{\text{grid}}=64$ emulator. The optimal hyperparameters are summarized in table 3. The best-performing emulators achieved $R^{2}$ scores of 0.969 and 0.990 on the test set, for $N_{\text{grid}}=32$ and 64, respectively. Additional diagnostics for the best-performing emulators, including the evolution of training and model-validation MSE and a scatter plot comparing true versus predicted $\chi^{2}$ values on the test set, are presented in figures 2 and 3.
@@
We incorporated the best-performing emulators described above into a conventional MCMC framework and used this new ANN-based framework to repeat the analyses described in section 2.3 for both $N_{\text{grid}}=32$ and 64. In figure 4 and 5, we show the comparison of the posteriors obtained using the conventional full MCMC at high-resolution (red) and the ANN-based MCMC (blue). The same priors were used for both cases. The figure demonstrates that the posterior distributions of the parameters obtained from both methods are in close agreement, establishing the reliability of the ANN-based MCMC framework.
@@
The results in table 4 constitute the primary validation of this framework. For the $N_{\text{grid}}=64$ case, the ANN-based approach yields approximately a $\sim 70$-fold reduction in total CPU hours while producing posterior constraints that are statistically indistinguishable from those obtained with the full forward model. This improvement is not merely incremental; it demonstrates that the method is capable of transforming the computational landscape of reionization inference. In particular, it establishes the feasibility of analyses that would otherwise be prohibitive, such as constraining the fourteen-parameter model [choudhuryCapturingSmallScaleReionization2025] required for JWST datasets, which remains computationally intractable using conventional MCMC techniques.
@@ 4 Summary and Future Outlook
We note that while this work focused on a feed-forward ANN, the field is rapidly adopting more advanced machine learning techniques, such as Simulation-Based Inference (SBI) using normalizing flows or diffusion models [e.g., alsingFastLikelihoodfreeCosmology2019, hassanHIFlowGeneratingDiverse2022, zhaoCanDiffusionModel2023]. These methods, however, still face the fundamental challenge of sampling efficiency. They perform optimally when training points are concentrated in or near the high-probability posterior region. Therefore, the core methodological contribution of this paper, using a reliable, coarse-resolution MCMC to efficiently identify the high-likelihood region, is a general-purpose strategy. It can be directly adapted to generate optimal training or proposal sets for these next-generation inference tools, significantly boosting their efficiency and making them more viable for complex, high-dimensional astrophysical problems.
This substantial gain in computational efficiency is not merely incremental, it is enabling. The five-parameter model explored here serves as a robust proof of concept, demonstrating that our emulator framework faithfully reproduces full high-resolution inference at a fraction of the cost. The most immediate and scientifically pressing application is to extend this approach to more complex models that are currently computationally prohibitive. In particular, our next step is to apply this methodology to the fourteen-parameter framework of ref. [choudhuryCapturingSmallScaleReionization2025], which is required to incorporate the latest JWST UV luminosity function measurements in a self-consistent manner. Exploring such a high-dimensional space is effectively impossible with conventional MCMC techniques, and the emulator strategy presented here provides a practical and scalable path forward.

### v2

@@ 2.3 Parameter constraints using conventional MCMC
For all observational constraints, except (2) and (3), we use a multidimensional Gaussian likelihood of the form
@@
where the index $i$ runs over all the data points used. For data with asymmetric error bars, we use the upper uncertainty when the model estimate lies above the data point, and the lower uncertainty when it lies below. For constraints (2) and (3), where we have limits instead of measurements, we use a likelihood given by [33, 53]
@@
To sample the parameter space and generate the posterior probability distributions, we use the publicly available package Cobaya [80, 81]. Specifically, we use its implementation of the Metropolis-Hastings sampler [59, 48]. The samples are drawn using 8 parallel chains [47, 48]. The chains are assumed to be converged when the Gelman-Rubin R1R-1 statistic [32] falls below 0.01 for two consecutive evaluations. Once converged, we discard the first 30% of the samples as burn-in and use the rest for further analysis. The resultant posterior distributions are analyzed using the GetDist [49] package. We use simulations with two different resolutions to constrain the free parameters, one with a grid size of $\Delta x=8~h^{-1}$ cMpc ($32^{3}$ grid cells), and another with a grid size of $\Delta x=4~h^{-1}$ cMpc ($64^{3}$ grid cells). The resultant parameter constraints for the two resolutions are shown in red in figures 5 and 6. The significant computational cost of this full MCMC run, which is quantified later (see table 4), provides the primary motivation for developing the efficient ANN-based emulator framework described in the next section.
@@ 3 ANN-based emulator
In this section, we present our ANN-based emulator approach, beginning with a brief overview of ANNs, followed by a detailed description of how they are used to construct emulators of $\chi^{2}(\theta)$, including dataset generation, training methodology, and integration within an MCMC framework. The same observational constraints and priors used in the previous section are adopted to enable a direct comparison between the two approaches. All the emulators discussed here are implemented using the TensorFlow Keras API [1, 13]. Figure 1 provides a schematic overview of the emulator-based inference framework’s methodology and workflow.
@@ 3.1 A brief on ANNs
by iteratively updating the weights and biases of the network over multiple training cycles, called epochs, with updates occurring after processing mini-batches of data. The training uses an algorithm called backpropagation [68], which applies the chain rule to compute gradients of the loss function with respect to the weights and biases. These gradients are then used by an optimization algorithm to update the network’s weights and biases. The step size of these updates is controlled by the learning rate, which determines how quickly the model adjusts its weights during training.
@@
Figure 1: Schematic overview of the emulator-based inference framework’s methodology and workflow.
@@ 3.2 General methodology for emulator training and MCMC integration
Building upon the basic principles outlined above, our primary objective is to train an ANN-based emulator that can accurately and rapidly interpolate the chi-squared value, $\chi^{2}(\theta)$, for any given parameter vector $\theta$ within the prior volume. In particular, the emulator must interpolate accurately in the high-likelihood (low $\chi^{2}$) region, where precise evaluations of $\chi^{2}(\theta)$ are most critical for parameter inference. This trained emulator can then be embedded within a conventional MCMC framework to efficiently explore the parameter space. This section outlines the general framework we use for training, validating, and deploying these emulators.
We choose to emulate $\chi^{2}(\theta)$ directly rather than the individual observables entering the likelihood. This approach significantly simplifies the learning task: $\chi^{2}$ is expected to be a smooth scalar function of the parameters, avoiding the challenges of jointly emulating several observables with different dynamic ranges and noise properties. This choice has also proven robust in earlier emulator frameworks for SCRIPT [55, 18], and it provides a scalable strategy for higher-dimensional extensions of the model. However, emulating the likelihood directly also has some limitations compared to emulating individual observables. In particular, the emulator is tied to a fixed observational dataset and likelihood definition, and would need to be retrained if either were modified. In addition, errors in the emulator predictions are not easily attributable to specific observables, reducing interpretability compared to observable-level emulators. Nevertheless, for the present work, these limitations are outweighed by the practical advantages of emulating the scalar $\chi^{2}(\theta)$. In particular, this choice allows us to use a considerably simpler ANN architecture and, consequently, achieve significantly faster and more stable training.
@@
The emulator is implemented as a multilayer perceptron (MLP) with Rectified Linear Unit (ReLU) activations in all hidden layers. Optimization is performed using the AdamW optimizer [50], with the weight decay factor fixed at its default value of 0.004, and all layer weights initialized using the Glorot scheme [35]. A given training dataset (described in section 3.3) is partitioned into training, model-validation, and test subsets in the ratio 6:2:2. The input parameter vectors are then standardized by subtracting, for each parameter, its mean value over the training set and rescaling by the corresponding standard deviation. These scaling parameters are computed using only the training set and then applied unchanged to the model-validation and test sets, ensuring that no information from the evaluation data is used during training. The batch size is chosen to scale with the size of the training set and is set to the integer part of its square root.
@@
During each training stage, a grid of candidate networks is trained on the training set, systematically varying the number of layers, the number of neurons per layer, and the learning rate, with the explored ranges listed in table 2. To mitigate the effect of unfavorable weight initialization, which can sometimes yield poor results even for otherwise well-chosen hyperparameters, each candidate network is trained twice using different random seeds. Training continues until the MSE on the model-validation set fails to decrease for 50 consecutive epochs, implementing early stopping to prevent overfitting. Once all candidate networks have been trained, they are evaluated on the test set, and the best-performing emulator is chosen as the trained network with the largest value of
@@ 3.3.1 Results using Latin Hypercube sampling
Figure 2: Parameter constraints from the full high-resolution ($N_{\text{grid}}=32$) MCMC run (red) and the ANN-based MCMC runs with the emulator trained on $10^{4}$ parameter vectors sampled using LH sampling from the full prior, and with the priors reduced by factors of 2 and 4.
@@ 3.3.2 Training set using targeted sampling
In this work, we instead follow the targeted sampling strategy introduced in [55], which takes advantage of two features of our setup: (i) the large-scale properties of our model are convergent with respect to simulation resolution, and (ii) full MCMC runs with coarse-resolution simulations are computationally inexpensive. This allows us to sample parameter vectors for the training dataset directly from the MCMC chain of a full MCMC run using inexpensive coarse-resolution simulations, ensuring that samples are concentrated in the high-likelihood regions of parameter space.
@@
In addition, we define a fixed validation set, $\{\theta\}_{\mathrm{val}}$, at the outset. This set is used exclusively to evaluate the convergence of the adaptive sampling procedure and is distinct from the “model-validation” set described in section 3.2, which is used during network training for early stopping. This set is constructed by drawing samples with probability proportional to their weights, ensuring that it provides an unweighted sample representative of the full posterior distribution. Crucially, this set is never used during training, as doing so would cause data leakage, where information from the validation set inadvertently influences training, resulting in overly optimistic performance estimates and poor generalization to unseen parameters. After each training stage, the emulator predicts $\chi^{2}$ for the validation set. The training dataset is then augmented with 300 additional samples, a new emulator is retrained, and the validation predictions are updated. To measure convergence, we compute the KL divergence between successive validation predictions,
@@ 3.4 Parameter constraints using ANN-based MCMC
For our setup, convergence of the adaptive sampling procedure required a training dataset size of 1200 samples for the $N_{\text{grid}}=32$ emulator, and 900 for the $N_{\text{grid}}=64$ emulator. The optimal hyperparameters are summarized in table 3. These hyperparameters were selected for optimization because they dominantly control the model’s expressive capacity and the convergence properties of the optimizer. Other hyperparameters (e.g., weight decay and batch size) were held fixed to limit the computational cost of the search and to avoid introducing degeneracies, particularly between batch size and learning rate. In addition, since early stopping already provides effective regularization, we fix the weight decay at its default value rather than tuning it explicitly. The best-performing emulators achieved $R^{2}$ scores of 0.969 and 0.990 on the test set, for $N_{\text{grid}}=32$ and 64, respectively. Additional diagnostics for the best-performing emulators, including the evolution of training and model-validation MSE and a scatter plot comparing true versus predicted $\chi^{2}$ values on the test set, are presented in figures 3 and 4.
@@
We incorporated the best-performing emulators described above into a conventional MCMC framework and used this new ANN-based framework to repeat the analyses described in section 2.3 for both $N_{\text{grid}}=32$ and 64. In figure 5 and 6, we show the comparison of the posteriors obtained using the conventional full MCMC at high-resolution (red) and the ANN-based MCMC (blue). The same priors were used for both cases. The figure demonstrates that the posterior distributions of the parameters obtained from both methods are in close agreement, establishing the reliability of the ANN-based MCMC framework.
@@
The results in table 4 constitute the primary validation of this framework. For the $N_{\text{grid}}=64$ case, the ANN-based approach yields approximately a $\sim 70$-fold reduction in total CPU hours while producing posterior constraints that are statistically indistinguishable from those obtained with the full forward model. This improvement is not merely incremental; it demonstrates that the method is capable of transforming the computational landscape of reionization inference. In particular, it establishes the feasibility of analyses that would otherwise be prohibitive, such as constraining the fourteen-parameter model [19] required for JWST datasets, which remains computationally intractable using conventional MCMC techniques.
@@ 4 Summary and future outlook
We note that while this work focused on a feed-forward ANN, the field is rapidly adopting more advanced machine learning techniques, such as Simulation-Based Inference (SBI) using normalizing flows or diffusion models [e.g., 2, 38, 85]. These methods, however, still face the fundamental challenge of sampling efficiency. They perform optimally when training points are concentrated in or near the high-probability posterior region. In this context, two methodological elements employed in this work are of direct relevance: the use of a reliable coarse-resolution MCMC to efficiently identify the high-likelihood region of parameter space, and the adaptive determination of the training dataset size based on emulator convergence. Together, these elements provide a general-purpose strategy for generating optimal training or proposal sets for these next-generation inference tools, significantly boosting their efficiency and making them more viable for complex, high-dimensional astrophysical problems.
This substantial gain in computational efficiency is not merely incremental, it is enabling. The five-parameter model explored here serves as a robust proof of concept, demonstrating that our emulator framework faithfully reproduces full high-resolution inference at a fraction of the cost. The most immediate and scientifically pressing application is to extend this approach to more complex models that are currently computationally prohibitive. In particular, our next step is to apply this methodology to the fourteen-parameter framework of ref. [19], which is required to incorporate the latest JWST UV luminosity function measurements in a self-consistent manner. Exploring such a high-dimensional space is effectively impossible with conventional MCMC techniques, and the emulator strategy presented here provides a practical and scalable path forward.